# Settling dynamics of non-Brownian suspension of spherical and cubic particles in Stokes flow

### v1

@@ Abstract
The present study investigates the gravity-driven settling dynamics of non-Brownian suspensions of spherical and cubic particles within a triply periodic domain. The effect of solid volume fraction on the evolving microstructure of a suspension is numerically examined by employing the Rigid MultiBlob method under Stokes flow conditions. Our simulations accurately reproduce macroscopic trends observed in experiments and show strong agreement with established semi-empirical correlations over a wide range of volume fractions. It is observed that the hydrodynamic interactions between particles and the surrounding fluid dominate the settling mechanism at low to moderate solid volume fractions, whereas, frequent collisions between particles in a highly packed space tend to suppress velocity fluctuations at denser regime. The transport properties for dilute suspensions are found to be primarily shaped by an anisotropic microstructure, though this anisotropy decreases as many-body interactions significantly grow at higher volume fractions. Notably, cubic particles exhibit lower anisotropy in velocity fluctuations compared to spheres, due to more effective momentum and energy transfer from the gravity-driven direction to transverse directions. Moreover, cubic particles also display higher velocity fluctuations and particle rebounding after collisions, which may create a transient tendency for cubes to move against gravity more readily than spheres. The distinctive behaviors of cubic and spherical particles can notably influence the efficiency of size segregation in bidisperse mixtures. Interestingly, local velocity fluctuation intensity increases for both particle types in mixed suspensions, affecting the local sedimentation velocity field. Overall, spherical particles are found to settle faster than cubes at low to moderate volume fractions.
@@ 2 Governing equations and numerical method
Figure 1: Schematic representation of the fully periodic cubic computational box consisting of a randomly placed spherical/cubic particles. The spherical body is discretized with a resolution of $42$ small blue blobs, whereas a resolution of $125$ blobs is considered for cubic particle.
The applied numerical set up is illustrated in Fig. 1. Two particles of different types are investigated namely: spherical (left) and cubic (right) particles. In simulations, monodisperse spherical/cubic particles are allowed to settle under the action of gravity in a quiescent viscous fluid under creeping flow conditions, i.e. $\text{Re}=0$. The suspension is composed with $M$ rigid particles denoted by $\{\Im_{p}\}_{p=1}^{M}$, where each particle is tracked by its position $\boldsymbol{q}_{p}$ and its orientation $\boldsymbol{\theta}_{p}$. The linear and angular velocities of particle $p$ are denoted by $\boldsymbol{u}_{p}$ and $\boldsymbol{\omega}_{p}$, respectively. The total external force applied to particle $p$ is $\boldsymbol{f}_{p}$, and the total torque $\boldsymbol{\tau}_{p}$. The velocity and pressure, $\boldsymbol{v}$ and $p$, of the fluid with viscosity $\eta$ are governed by the Stokes equations
@@
In the Stokes limit the conservation of linear and angular momentum reduces to the balance of force and torque and thus, the hydrodynamic traction, $-\boldsymbol{\lambda}$, integrated on the particles surface balance the external forces and torques applied on them
@@
We construct rigid bodies of essentially arbitrary shape as a collection of rigidly connected ”blobs”, each having a finite radius $a$ and a position $\boldsymbol{r}_{i}$, (see Fig. 1), forming a composite object that we will refer to as a Rigid MultiBlob (RMB). The hydrodynamic problem is solved by RMB method, which is a regularized boundary integral method, formulated as a symmetric saddle system involving the mobility matrix between markers. A detailed description of this method can be found in the study of Kallemov et al. [38] and Balboa et al. [39]. After discretizing the particles, the integrals in the balance of force and torque (eq. (4)) become sums over the blobs, while the no-slip condition is applied on each blob
@@
This regularization of the Green’s function makes the mobility $\boldsymbol{M}$ positive definite and the numerical method quite robust and easy to implement. To include the effect of the PBC, we compute the action of the mobility matrix with a Fast Multipole Method that accounts for the periodic images with a linear computational cost [42].
@@
where $\boldsymbol{K}$ is a block matrix that converts the particles velocities into blobs velocities and its transpose sums over the force on the blobs to give the total force-torque acting on the particles. For a many particle suspension (7) is a large linear system than can be solved efficiently by a preconditioned GMRES iterative solver [39]. Finally, the spheres and cubes are pulled by an external force along the $z$-axis to model the pull of gravity. Also, an additional repulsive force is also incorporated to account for the surface roughness or steric interactions between solid particles, effectively preventing artificial overlaps. The force, acting between blobs forming the colloids, is derived from the potential
@@ 3 Results and Discussion
Symbols Description          Unit
$M$ No. of particles         $125$
$r_{s}$ Radius of each blob for sphere    $0.2435\,$\mu\mathrm{m}$$
$r_{c}$ Radius of each blob for cube    $0.1091\,$\mu\mathrm{m}$$
@@
[…]
$R_{s}$ Radius of sphere        $1\,$\mu\mathrm{m}$$
$R_{c}$ Side length of cube        $1\,$\mu\mathrm{m}$$
$\psi_{0}$ Repulsion strength        $0.0957\,$\mathrm{p}\mathrm{N}\mu\mathrm{m}$$
$b$ Decay length spheres        $0.02435\,$\mu\mathrm{m}$$
$b$ Decay length cubes        $0.01091\,$\mu\mathrm{m}$$
Three-dimensional simulations are performed within a cubic, triply periodic domain, illustrated in Fig. 1. The computational domain has dimensions of $L\times L\times L$, where $L$ is the length of each side. To achieve the desired solid volume fraction, $\phi$, we fix the total number of particles, $M$, and adjust the size of the computational box. The solid volume fraction is then calculated as $\phi=MV/L^{3}$, where $V$ represents the volume of each particle, either being spherical or cubic in shape. The characteristic length scale is indicated by $R$, chosen as the radius, $R_{s}$, for spheres (side length, $R_{c}$, for cubes). We use dynamic simulations to obtain the trajectory and velocities of sedimenting particles. Particles are initially seeded at random positions ensuring that particles do not overlap. Then, the trajectories are computed for a finite time and ensemble averages are formed over all the particles as time goes on. In our coordinate system, gravity is aligned in the negative $z$-direction, also referred to as vertical direction in the following. The horizontal direction corresponds to the $x$ and $y$-coordinates. The physical parameters and computational settings used in the present study is listed in Table LABEL:table:Table_1.
@@ 3.1 Sedimentation of two spheres
We begin our investigation by validating the steady motion of a pair of identical and equal sized spheres settling under gravity in unbounded domain (i.e. $L\rightarrow\infty$). The initial configuration of two spheres is shown in Fig. 2a. The spheres are either separated horizontally or vertically to the line of center. As pointed out by Batchelor [5], the settling velocity depends on the separation between the particles and the orientation of the line of centers to the vertical. Later, Maxey and Patel [43] introduced a finite-valued force multipole expansion to describe the dynamics of spherical particles sedimenting in Stokes flow. They used the Force Coupling Model (FCM) to evaluate the settling velocity of two equal spheres in the above mentioned configurations. If the particles are otherwise isolated and not influenced by other particles or boundaries, they will have equal settling velocities due to the symmetry of the induced Stokes flow. The mutual influence of the disturbance flow created by each particle leads to a settling velocity larger than that for a single isolated particle. Our numerical result in Fig. 2b shows a good agreement with the result of Maxey and Patel [43]. The settling velocity $U_{\text{sed}}$ is scaled by the Stokes velocity $U_{0}$ of a single sphere and given in terms of the distance between the centers, so $r/R=2$ corresponds to touching spheres. In general there is no external torque on the particles and they are free to rotate in response to the local fluid velocity. Fig. 2b illustrates that the particles settle with the same velocity as an isolated sphere in an infinite medium, when they are very far apart from each other (i.e., $r/R\to\infty$). However, the mutual influence of the flow disturbance created by each particle results in a settling velocity that is larger than that of a single isolated particle. The no-slip condition between the spheres prevents the rotation of the spheres at or near contact leading to a sharp, local change in settling velocity with $U_{\text{sed}}/U_{0}\approx 1.41$ at $r/R=2.01$.
@@ 3.2 Sedimentation velocity of monodisperse suspension of spheres and cubes
Figure 4: Average settling velocity $U_{\text{sed}}$ of the particles normalized by the Stokes velocity $U_{0}$ as a function of volume fraction ($\phi$) to highlight the hindered settling of the particles. Solid lines denote our simulation result; dashed lines show various theoretical predictions of Hayakawa and Ichiki [46], Richardson and Zaki [47], Di Felice [9], Batchelor [5], simulation result of Yamamoto et al. [48], and the symbols represent the experimental data of Nicolai et al. [14].
@@ 3.3 Velocity fluctuations
During settling, particle velocities fluctuate both vertically and horizontally due to inter-particle hydrodynamic interactions. Vertical velocity fluctuations reflect temporal deviations from the mean sedimentation velocity, while horizontal fluctuations, occurring perpendicular to gravity, are characterized by the mean velocities in the $XY$ plane. The time-averaged standard deviations of particle velocities parallel and perpendicular to gravity are quantified as $\Delta U_{z}=\sqrt{\langle[U_{iz}-U_{\text{sed}}]^{2}\rangle}$ and $\Delta U_{xy}=\sqrt{\langle U_{ix}^{2}+U_{iy}^{2}\rangle}$, respectively. Here, $U_{iz}$ represents the temporal velocity of the $i$-th particle along the gravitational direction, while $U_{ix}$ and $U_{iy}$ are the velocities in the horizontal directions. These standard deviation values are normalized by the corresponding settling velocity of each suspension, commonly referred to as ’relative fluctuation’ [49, 14]. Figures 5a and 5b display dimensionless velocity fluctuations of particles in the settling direction ($\sigma_{\parallel}=\Delta U_{z}/U_{\text{sed}}$) and in the perpendicular direction ($\sigma_{\perp}=\Delta U_{xy}/U_{\text{sed}}$), respectively. The fluctuations exhibit similar qualitative behavior in both directions. At low volume fractions ($\phi<0.04$), $\sigma_{\parallel}$ and $\sigma_{\perp}$ display characteristics of the Stokes regime, with a $\phi^{1/2}$ dependence (see insets of Figs. 5a,b). As the solid volume fraction increases, velocity fluctuations rise gradually up to a moderate range ($0.04\leq\phi\leq 0.2$). Beyond this range ($\phi>0.2$), a gradual decline in fluctuations is observed for both directions. Additionally, velocity fluctuations are slightly higher for cubes compared to spheres, with the differences between the two shapes becoming more pronounced as $\phi$ increases. As the volume fraction increases within the range ($0.04\leq\phi\leq 0.2$), particle-particle interactions driven by collisions become prominent, leading to an increase in velocity fluctuations due to intensified interparticle hydrodynamic interactions. Consequently, momentum is significantly transferred from the axial direction to the transverse direction, which amplifies velocity fluctuations in the transverse direction [50]. Notably, even a slight reorientation of a cube can generate considerable hydrodynamic torques, leading to rotations that further enhance this effect [36]. As a consequence, cube seems to experience higher transverse velocity fluctuations compared to spheres, thus facilitating a more efficient transfer of momentum from the axial to the transverse direction. Further increment in $\phi$, the average inter-particle distance, which is inversely proportional to ($\phi$), diminishes, causing particles to become more densely packed [51]. Beyond ($\phi>0.2$), this dense packing restricts particle motion, resulting in a reduction in velocity fluctuations. In this range, the reduced inter-particle spacing disrupts the fluid structure required for particle clustering, while the frequency of hydrodynamic collisions rises, further suppressing velocity fluctuations.
@@
The amplification in velocity fluctuations with increasing solid volume fraction is closely linked to the degree of anisotropy in fluctuations for both cubic and spherical suspensions. This anisotropy, defined as the ratio of vertical to horizontal velocity fluctuations ($\Gamma=\sigma_{\parallel}/\sigma_{\perp}$), is shown in Fig. 5c. In the most dilute cases, anisotropy is notably high for both particle shapes, primarily due to the restricted lateral movement in this regime. For dilute suspensions ($\phi\leq 0.05$), $\Gamma$ ranges between 1.7 and 2.1, consistent with observations from the experimental work of Nicolai and Guazzelli [15] on non-Brownian spheres in very dilute conditions. As the solid volume fraction increases to moderate levels ($\phi\approx 0.2$), velocity fluctuations rise in both vertical and horizontal directions, leading to a significant reduction in anisotropy. In this range, anisotropy decreases by up to $30\%$ compared to the dilute regime, with $\Gamma$ approaching approximately $1.22$. A similar trend was also noted in the studies of Nicolai and Guazzelli [15] and Ladd [16]. Notably, the anisotropy in cubic suspensions is consistently lower than in spherical suspensions. As the volume fraction continues to increase, the level of anisotropy further diminishes due to the more frequent and multidirectional inter-particle collisions.

### v2

@@ Abstract
The present study investigates the gravity-driven settling dynamics of non-Brownian suspensions consisting of spherical and cubic particles within a triply periodic domain. We numerically examine the impact of solid volume fraction on the evolving microstructure of the suspension using the Rigid MultiBlob method under Stokes flow conditions. Our simulations match macroscopic trends observed in experiments and align well with established semi-empirical correlations across a broad range of volume fractions. At low to moderate solid volume fractions, the settling mechanism is primarily governed by hydrodynamic interactions between the particles and the surrounding fluid. However, frequent collisions between particles in a highly packed space tend to suppress velocity fluctuations at denser regimes. For dilute suspensions, transport properties are predominantly shaped by an anisotropic microstructure; though, this anisotropy diminishes as many-body interactions intensify at higher volume fractions. Notably, cubic particles exhibit lower anisotropy in velocity fluctuations compared to spherical particles, owing to more efficient momentum and energy transfer from the gravity-driven direction to transverse directions. Finally, bidisperse suspensions with mixed particle shapes show enhanced velocity fluctuations, driven by shape-induced variations in drag and increased hydrodynamic disturbances. These fluctuations in turn affect the local sedimentation velocity field, leading to the segregation of particles in the mixture.
@@ 2 Governing equations and numerical method
Figure 1: Schematic representation of the fully periodic cubic computational box consisting of a randomly placed spherical/cubic particles. The spherical body is discretized with a resolution of $42$ small blobs, whereas a resolution of $125$ blobs is considered for cubic particle.
The applied numerical set up is illustrated in Fig. 1. Two particles of different types are investigated namely: spherical (left) and cubic (right) particles. In simulations, monodisperse spherical/cubic particles are allowed to settle under the action of gravity in a quiescent viscous fluid under creeping flow conditions, i.e. $\text{Re}=0$. The suspension is composed with $M$ rigid particles denoted by $\{\Im_{p}\}_{p=1}^{M}$, where each particle is tracked by the position of its center $\boldsymbol{q}_{p}$ and its orientation $\boldsymbol{\theta}_{p}$. We use unit quaternions to represent orientations but other formulations are formally equivalent [38]. The linear and angular velocities of particle $p$ are denoted by $\boldsymbol{u}_{p}$ and $\boldsymbol{\omega}_{p}$, respectively. The total external force applied to particle $p$ is $\boldsymbol{f}_{p}$, and the total torque around the center of the particle $\boldsymbol{\tau}_{p}$. The velocity and pressure, $\boldsymbol{v}$ and $p$, of the fluid with viscosity $\eta$ are governed by the Stokes equations
@@
The equations are closed by the balance of force and torque. The flow exerts a traction on the particles surface $-\boldsymbol{\lambda}=\boldsymbol{\sigma}\cdot\boldsymbol{n}$ where $\boldsymbol{n}$ is the surface normal and $\boldsymbol{\sigma}=-p\boldsymbol{I}+\eta\left(\boldsymbol{\nabla}\boldsymbol{v}+(\boldsymbol{\nabla}\boldsymbol{v})^{T}\right)$ the fluid stress. This hydrodynamic traction, integrated on the particles surfaces, balance the external forces and torques applied to the particles
@@
We construct rigid bodies of essentially arbitrary shape as a collection of rigidly connected ”blobs”, each having a finite radius $a$ and a position $\boldsymbol{r}_{i}$, (see Fig. 1), forming a composite object that we will refer to as a Rigid MultiBlob (RMB). The hydrodynamic problem is solved by RMB method, which is a regularized boundary integral method, formulated as a symmetric saddle system involving the mobility matrix between markers. A detailed description of this method can be found in Refs. [39, 40]. After discretizing the particles, the integrals in the balance of force and torque (eq. (4)) become sums over the blobs, while the no-slip condition is applied on each blob
@@
This regularization of the Green’s function makes the mobility $\boldsymbol{M}$ positive definite and the numerical method quite robust and easy to implement. To include the effect of the periodic boundary condition (PBC), we compute the action of the mobility matrix with a Fast Multipole Method that accounts for the periodic images with a linear computational cost [43].
@@
where $\boldsymbol{K}$ is a block matrix that converts the particles velocities into blobs velocities and its transpose sums over the force on the blobs to give the total force-torque acting on the particles. For a many particle suspension (7) is a large linear system than can be solved efficiently by a preconditioned GMRES iterative solver [40]. Finally, the spheres and cubes are pulled by an external force along the $z$-axis to model the pull of gravity. With the finite resolution used to discretize particles, see Fig. 1, lubrication forces are not recovered exactly, therefore, an additional repulsive force is incorporated to account for the steric interactions between solid particles, effectively preventing artificial overlaps. The force, acting between blobs forming the colloids, is derived from the potential
@@ 3 Results and Discussion
Symbol Description Value Comments
$M$ No. of particles $125$ unless otherwise state
$R$ Particle radius $1$ for cubes the volume-equivalent sphere
radius is used, the cube side is
$h=(4\pi/3)^{1/3}R$
$r_{s}$ Blob radius for sphere $0.2435$ spheres are discretized with 42 blobs
$r_{c}$ Blob radius for cube $0.1746$ cubes are discretized with 125 blobs
@@
$\eta$ Fluid viscosity $1$
$f$ Gravity force per particle $1$
$\psi_{0}$ Repulsion strength $0.011$ for spheres and cubes
$b$ Decay length spheres $r_{s}/10$
$b$ Decay length cubes $r_{c}/10$
Three-dimensional simulations are performed within a cubic, triply periodic domain, illustrated in Fig. 1. The computational domain has dimensions of $L\times L\times L$, where $L$ is the length of each side. To achieve the desired solid volume fraction, $\phi$, we fix the total number of particles, $M$, and adjust the size of the computational box. The solid volume fraction is then calculated as $\phi=MV/L^{3}$, where $V$ represents the volume of each particle, either being spherical or cubic in shape. The characteristic length scale, denoted by $R$, is defined as the radius for spherical particles and the volume-equivalent sphere radius for cubes, i.e. the cubes have side length $h=\left(4\pi/3\right)^{1/3}R$. We use dynamic simulations to obtain the trajectory and velocities of sedimenting particles. Particles are initially seeded at random positions ensuring that particles do not overlap. Then, the trajectories are computed for a finite time and ensemble averages are formed over all the particles as time goes on. In our coordinate system, gravity is aligned in the negative $z$-direction, also referred to as vertical direction in the following. The horizontal direction corresponds to the $x$ and $y$-coordinates. The physical parameters and computational settings used in the present study is listed in Table LABEL:table:Table_1.
@@ 3.1 Sedimentation of two spheres
We begin our investigation by validating the steady motion of a pair of identical and equal sized spheres settling under gravity in unbounded domain (i.e. $L\rightarrow\infty$). The initial configuration of two spheres is shown in Fig. 2a. The spheres are either separated horizontally or vertically to the line of center. As pointed out by Batchelor [5], the settling velocity depends on the separation between the particles and the orientation of the line of centers to the vertical. Later, Maxey and Patel [44] introduced a finite-valued force multipole expansion to describe the dynamics of spherical particles sedimenting in Stokes flow. They used the Force Coupling Model (FCM) to evaluate the settling velocity of two equal spheres in the above mentioned configurations. If the particles are otherwise isolated and not influenced by other particles or boundaries, they will have equal settling velocities due to the symmetry of the induced Stokes flow. The mutual influence of the disturbance flow created by each particle leads to a settling velocity larger than that for a single isolated particle. Our numerical result in Fig. 2b shows a good agreement with the result of Maxey and Patel [44]. The settling velocity $U_{\text{sed}}$ is scaled by the Stokes velocity $U_{0}$ of a single sphere and given in terms of the distance between the centers, so $r/R=2$ corresponds to touching spheres. There is no external torque on the particles and they are free to rotate in response to the local fluid velocity. Fig. 2b suggests that the particles settle with the same velocity as an isolated sphere in an infinite medium, when they are very far apart from each other (i.e., $r/R\to\infty$). However, the mutual influence of the flow disturbance created by each particle results in a settling velocity that is larger than that of a single isolated particle. This effect leads to a velocity ratio of $U_{\text{sed}}/U_{0}\approx 1.41$ at $r/R=2.01$ for horizontally separated spheres.
3.2 Effect of domain size
Figure 3: Average settling velocity of spherical particles versus the number of particles $M$ at two different solid volume fractions: (a) $\phi=0.025$ and (b) $\phi=0.36$. The system size $L$ is changed according with $M$ to keep the volume fraction constant, the settling velocity is normalized by $U_{0}$, and error bars represent the standard deviation of the mean settling velocity. The red symbols represent the value $M=125$ which we use in the rest of our simulations.
The system size determines how well the system approximates an infinite suspension and whether artificial effects distort the settling dynamics [45, 46]. To ensure the system size is sufficiently large we study its influence on the average sedimenting velocity of spherical particles in two cases: a dilute suspension ($\phi=0.025$) and a moderately dense suspension ($\phi=0.36$). In both cases we vary the number of particles, $M$, and the system size, $L$, while keeping the volume fraction constant and measure the average sedimenting velocity. The results indicate that in both cases, a small box size significantly affects the mean settling velocity, see Figs. 3a,b. However, as the domain size increases it soon reaches the asymptotic value. In the rest of the work we use $M=125$ so the finite sizes effects are small while the computational cost is still moderate. The red symbols in Figs. 3a,b indicate the selected box size used in our simulations for different solid volume fractions.
@@ 3.3 Sedimentation velocity of monodisperse suspension of spheres and cubes
Figure 5: Average settling velocity $U_{\text{sed}}$ of the particles normalized by the Stokes velocity $U_{0}$ as a function of volume fraction ($\phi$) to highlight the hindered settling of the particles. Solid lines denote our simulation result; dashed lines show various theoretical predictions of Hayakawa and Ichiki [49], Richardson and Zaki [50], Di Felice [9], Batchelor [5], simulation result of Yamamoto et al. [51], and the symbols represent the experimental data of Nicolai et al. [14]. Here, the error bars are smaller than the symbols’ size.
@@ 3.4 Velocity fluctuations
During settling, particle velocities fluctuate both vertically and horizontally due to inter-particle hydrodynamic interactions. Vertical velocity fluctuations reflect temporal deviations from the mean sedimentation velocity, while horizontal fluctuations, occurring perpendicular to gravity, are characterized by the mean velocities in the $XY$ plane. The time-averaged standard deviations of particle velocities parallel and perpendicular to gravity are quantified as $\Delta U_{z}=\sqrt{\langle[U_{iz}-U_{\text{sed}}]^{2}\rangle}$ and $\Delta U_{xy}=\sqrt{\langle U_{ix}^{2}+U_{iy}^{2}\rangle}$, respectively. Here, $U_{iz}$ represents the temporal velocity of the $i$-th particle along the gravitational direction, while $U_{ix}$ and $U_{iy}$ are the velocities in the horizontal directions. These standard deviation values are normalized by the corresponding settling velocity of each suspension, commonly referred to as ’relative fluctuation’ [45, 14]. Figures 6a and 6b display dimensionless velocity fluctuations of particles in the settling direction ($\sigma_{\parallel}=\Delta U_{z}/U_{\text{sed}}$) and in the perpendicular direction ($\sigma_{\perp}=\Delta U_{xy}/U_{\text{sed}}$), respectively. The fluctuations exhibit similar qualitative behavior in both directions. At low volume fractions ($\phi<0.04$), $\sigma_{\parallel}$ and $\sigma_{\perp}$ display characteristics of the Stokes regime, with a $\phi^{1/2}$ dependence (see insets of Figs. 6a,b). As the solid volume fraction increases, velocity fluctuations rise gradually up to a moderate range ($0.04\leq\phi\leq 0.2$). Beyond this range ($\phi>0.2$), a gradual decline in fluctuations is observed for both directions. Additionally, velocity fluctuations are slightly higher for cubes compared to spheres, with the differences between the two shapes becoming more pronounced as $\phi$ increases. As the volume fraction increases within the range ($0.04\leq\phi\leq 0.2$), particles are more crowded and collision events, where steric interactions push the particles apart, become more common. Consequently, momentum is significantly transferred from the axial direction to the transverse direction, which amplifies velocity fluctuations in the transverse direction [52]. Notably, even a slight reorientation of a cube can generate considerable hydrodynamic torques, leading to rotations that further enhance this effect [36]. As a consequence, cube seems to experience higher transverse velocity fluctuations compared to spheres, thus facilitating a more efficient transfer of momentum from the axial to the transverse direction. Further increment in $\phi$, the average inter-particle distance, which is inversely proportional to ($\phi$), diminishes, causing particles to become more densely packed [53]. Beyond ($\phi>0.2$), this dense packing restricts particle motion, resulting in a reduction in velocity fluctuations. In this range, the reduced inter-particle spacing disrupts the fluid structure required for particle clustering, while the frequency of hydrodynamic collisions rises, further suppressing velocity fluctuations.
@@
The increase in velocity fluctuations with rising solid volume fraction is closely associated with the degree of anisotropy in these fluctuations for both cubic and spherical suspensions. This anisotropy, defined as the ratio of vertical to horizontal velocity fluctuations ($\Gamma=\sigma_{\parallel}/\sigma_{\perp}$), is shown in Fig. 6c. In highly dilute conditions, anisotropy is notably high for both particle shapes, primarily due to restricted lateral movement in this regime. For dilute suspensions ($\phi\leq 0.05$), $\Gamma$ ranges between 1.7 and 2.1 for both cases, aligning with the experimental findings of Nicolai and Guazzelli [15] on non-Brownian spheres under very dilute conditions. As the solid volume fraction increases to moderate levels ($\phi\approx 0.2$), velocity fluctuations intensify in both vertical and horizontal directions, leading to a substantial reduction in anisotropy. In this range, the anisotropy decreases compared to the dilute regime, with $\Gamma$ approaching approximately $1.3$ for cubic suspensions. A similar trend was also observed in the studies of Nicolai and Guazzelli [15] and Ladd [16]. In a sedimentation study at finite Reynolds numbers by Seyed-Ahmadi and Wachs [37] was observed that the higher rotational velocities of cubes contributed to the transfer of momentum and energy from the gravity to the transverse direction, which led to a lower velocity fluctuations anisotropy in cubic suspensions than in spherical suspensions. Here, we observe the same effect in zero Reynolds number regime. As $\phi$ increases further to $0.5$, the anisotropy continues to decline due to more frequent and multidirectional inter-particle collisions.